OpenStudy (anonymous):

Did I do this Geometry problem correctly? write an equation for a circle with a diameter that has endpoints at (-4 -1) and (-8 -9) Round to the nearest 10th if necessary. I got (x+6)^2+(y+5)^2=80

hero (hero):

What was the radius of this circle?

OpenStudy (anonymous):

All the information that's given, I stated.

hero (hero):

You didn't answer my question. I asked you if you knew the radius of the circle. You should know since you computed it. Otherwise, you wouldn't have found r^2 = 80. So I assume you also know what r is. So please tell me what the radius is.

OpenStudy (anonymous):

Oh, okay. :)

OpenStudy (anonymous):

I believe it's 40

hero (hero):

And how exactly did you get 40? What steps did you perform to get that?

hero (hero):

Even if the radius was 40, you'd still have to explain how you got the 80. the formula includes r^2 which is radius squared.

OpenStudy (anonymous):

english is my third language, i cant really articulate that explanation. :

OpenStudy (anonymous):

especially given the time it is at the moment. HA HA :)

hero (hero):

Time doesn't matter. The only thing that matters is NOW. NOW is an in the moment thing that actually transcends time. NOW is something that cannot be measured or quantified.

hero (hero):

But anyway, back to the point. How did you get 40 and how did you get 80? Try your best to explain. Show some algebraic steps at least.

OpenStudy (anonymous):

When I say time, i mean that it's past the hour of which i would normally be asleep.

OpenStudy (anonymous):

8, 5. sorry... REALLY cant explain. could you please tell me if im right or wrong?

OpenStudy (anonymous):

maybe it's (x-6)^2 + (y-5)^2=20...

hero (hero):

Actually (-6,-5) is correct. And you seem to have the right radius so (x + 6)^2 + (y + 5)^2 = 20 is correct. By the way, if you are working on a problem and you come to a result, you need to make sure you do two things: 1. Be confident in your results 2. Be prepared to defend your answer In other words, you need to be sure of yourself when working with math.

hero (hero):

The reason for this is, if someone comes along and says "you are wrong", you can defend yourself by explaining why you are right.

hero (hero):

You will never be good at math unless you possess this ability.

OpenStudy (anonymous):

sorry, due to past expieriences with abuse and being degraded, i am never fully confident with myself.

hero (hero):

I'm sorry to hear that. The only thing that can save you is a mathematics video game.

OpenStudy (anonymous):

are you european

hero (hero):

Why does that matter? And why do you ask me if I'm European?

OpenStudy (anonymous):

to see if i can explain this to you in my native language. or a language im more comfortable speaking

OpenStudy (anonymous):

do you speak anything other than english?

hero (hero):

I'm American bro, sorry.

OpenStudy (anonymous):

Okay. Oregonian?